Problem 3.8 Find and sketch g(t)=x(t)*x(t) for the pair of functions shown below. Did you notice anything interesting? x(t) xxi(t) xz(t) xi(t) *xz(t) Note: Everlasting signal AAA xi(t) xalt) xi(t) + x2(t) Note: Everlasting signal АААА
Answer #1

x₂ (+) a) _ ^ vices B 22 4 6 + 24 X (+-7) Whem -ict co- - 6+t - 4+t -5 h When t <-1 -6+t -htt -5 h = (x₂ (T) x, (t-T) dt -6ttX, Lt) = u(t+2) xz(t) = ult+2)- ult-1) -2 - 2 1 Note ult) * u(t) = r(t) ultra) * u(t+b) = r(ta+6) X, (t) * x₂ (t) = u(+2) * (c) 1 X2 CT x (t) X (t-) an 6 2 T -2++ -2xt t Whenort <2 ct SE- +9dt When to O T -2t+1 z 2 Ty(t) = 0 + -2+t t 2 Jo When t > 1(7) ix 40 10) * (T) X2 (-T) x₂(t-T) yet) = ( x, (2). 4, (tt) de yee Yuce) de y = 0 4t to 4 t=h Area =0 Area=0 TyH) = 0) AsX, (t) * (t-1) -8tt yet) = 1 (0) 3 (4-7) de tot 1x,ce) dt Area=0 (y (+) =0 4 ) Ay [ Area=o 7 e) AM x, (t) - Xz(t) x2 Ct -T) 1